# Mapping Dark Matter Through the Dust of the Milky Way Part I: Dust Correction and Phase Space Density

### v1

@@ 3.3 Regularization
This regularization term penalizes solutions where $\epsilon\neq 1$ over large portions of the sky and diverges as $\epsilon$ approaches 00, when $f_{\rm corr}(\vec{x})$ is no longer defined. This term introduces an additional hyperparameter $\lambda_{\epsilon}$. A hyperparameter scan over $\lambda_{\epsilon}$ reveals that small values of $\lambda_{\epsilon}$ fail to set a scale for $\epsilon$, leading to potentially inconsistent scales between different trainings of $\epsilon$ which will later be averaged together. Very large values of $\lambda_{\epsilon}$ force $\epsilon=1$ across the entire sky, a solution that fails to correct any of the effects of dust. Through numerical experiments, we found that the range $\lambda_{\epsilon}\in[10^{-2},1]$ yields stable results, and we ultimately settled on $\lambda_{\epsilon}=10^{-1}$ as the canonical value for this analysis.
@@ 4 Results
In this section, we present the results of the trained MAF model of $f_{\rm obs}$ as well as the $\epsilon$ network. In a companion paper Ref. [43], we analyze the associated dust-corrected $\Phi$ model, where we examine the resulting acceleration and mass density field of the local Milky Way.
@@ 4.1 Dust Efficiency
    $\epsilon_{\rm L22}$ is computed from a Gaia dust map in the monochromatic band at $5500$, whereas the RC/RGB sample of stars with 6D kinematic information were observed in the narrow $\sim 8600$ band of the Gaia RVS. Depending on their chemical composition, dust clouds will have slightly different morphology when observed in different bands.
@@
Figure 3: Left: CBE-derived $\epsilon$ projected in Galactic coordinates at distances $d=1$, $2$, and $3$ kpc. At larger distances, familiar features in local dust clouds become more resolved in regions where $\epsilon<1$. However, not every feature seen here should be attributed to dust extinction. For example, the dark patch that first appears at $(\ell,b,d)\approx(5^{\circ},-4^{\circ},2\>\textrm{kpc})$ is an incomplete patch in the Gaia dataset due to crowding and saturation effects. Right: Same as left, using $\epsilon_{\rm L22}$ derived from the L22 dust map [42]. At $3$ kpc, $\epsilon_{\rm L22}$ approaches $1$ and is no longer reliable, as it cannot restore stars dimmed beyond the limits of the Gaia spectrometer.
@@
[…]
Figure 4: Left: Dust efficiency function $\epsilon$ estimated from the CBE along slices in the $x-y$ plane at $z=-0.25$ (top), $0.0$ (center), and $0.25$ kpc (bottom). The center row is a slice through the midplane of the Milky Way’s disk, where the most efficiency loss is expected due to the dense network of dust clouds. Right: $\epsilon_{\textrm{L22}}$ estimated from the extinctions provided by the L22 dust map, limited to within $d=2.5$ kpc (black circle). Beyond $d=2.5$ kpc, $\epsilon_{\rm L22}$ cannot restore missing stars and returns to $1$.
Figure 5: Comparison of $\epsilon_{\rm L22}$ derived from Ref. [42] (blue) and the $\epsilon$ estimated in this work (gold) along lines of sight centered on known significant dust clouds shown in Figure 1. The center line represents an average over a cone of opening angle $\theta=10$ arcmin. The error band for the $\epsilon$ model (blue) is dominated by measurement and statistical uncertainties. Likewise, the error band of the $\epsilon_{\rm L22}$ model (gold) includes the L22 dust map errors as well as an estimate of measurement and statistical errors. Both error bands include the subdominant variation across each circular section of the cone. The light gray band in some panels indicates lines of sight where $\epsilon_{\rm L22}$ degrades closer than $d=2.5$ kpc due to heavy dust extinction.
In Figure 5, we next examine $\epsilon$ evaluated within cones of radius $10$ arcmin that pass through the sixteen dusty regions of the Solar neighborhood highlighted in Figure 1. The uncertainties in $\epsilon$ shown in Figure 5 are derived by repeatedly retraining the MAFs and solving the CBE using multiple measurement error-smeared or resampled copies of the Gaia data, and by computing the variation across hundreds of retrainings with different network initializations. A full description of the uncertainty quantification procedure is given in Appendix C. The error band for $\epsilon_{\rm L22}$ is similarly derived, using the provided uncertainties in the L22 dust map [42] as well as error-smeared and resampled copies of the Gaia data. Additionally, for both $\epsilon$ and $\epsilon_{\rm L22}$, we take the average and variance over a short angular patch or “cone” in order to mitigate differences in resolution between the models. Variance across the cone is found to be subdominant to the primary measurement and statistical uncertainties.
@@
Figure 6: Pointwise “voxel” comparison between $\epsilon_{\rm L22}$ derived from Ref. [42] and the $\epsilon$ model that solves Eq. (1.3). Each efficiency value is evaluated over a grid of 3D voxels of volume $(25\>{\rm pc})^{3}$ that tile the space within $|z|<0.5$ kpc of and $d<2.5$ kpc. Top: Frequency of efficiency values in each voxel. Color scale indicates the density of voxels in each bin in a logarithmic scale. Bottom: Frequency of $\tau_{\epsilon}$ values between $\epsilon$ and $\epsilon_{\rm L22}$ accounting for errors. Dark blue indicates the $1-99\%$ percentile range of pull values. The solid black line indicates the average pull value.
As a final test, we compare $\epsilon$ and $\epsilon_{\rm L22}$ pointwise across a uniform, three-dimensional grid of nearby voxels in the Milky Way’s disk ($d<2.5$ kpc and $|z|<0.5$ kpc). In the left panel of Figure 6, we find that the models agree when there is no predicted extinction ($\epsilon\approx 1$). Additionally, we observe that $\epsilon$ lags behind as $\epsilon_{\rm L22}$ decreases from 1. This lag supports our previous observations that $\epsilon$ decreases from $1$ slightly later than $\epsilon_{\rm L22}$ along a given line of sight. This disagreement is slight, on the order of $||\epsilon-\epsilon_{L22}||\approx 0.1$ when both $\epsilon$ and $\epsilon_{\rm L22}>0.4$. In practice, at a given $\vec{x}$ within a region of moderate dust extinction, this means that these models on average disagree at the 10% level on the fraction of missing stars. This is consistent with the level of disagreement seen in Figure 5. In the right panel of Figure 6, we present the pull $\tau_{\epsilon}=(\epsilon-\epsilon_{\rm L22})/\sqrt{\sigma_{\epsilon}^{2}+\sigma_{\epsilon_{\rm L22}}^{2}}$ between $\epsilon$ and $\epsilon_{\rm L22}$ evaluated at each voxel when accounting for uncertainties. As can be seen from the pull distribution, the bulk of $\tau_{\epsilon}$ values skew slightly positive, indicating that the amount of dust extinction tends to be slightly underestimated relative to the L22 dust map. However, the absolute value of this pull rarely exceeds $1\sigma$, indicating excellent agreement between these two models.
@@ 4.2 Dust-Corrected PSD
Figure 7: Left: MAF model of the observed number density $n_{\textrm{obs}}$ of Gaia RC/RGB stars projected in Galactic coordinates at distances $d=1$ (top), $2$ (center), and $3$ kpc (bottom). Right: Extinction-corrected number density $n_{\rm corr}\equiv n_{\textrm{obs}}/\epsilon$.
Figure 8: Left: Histogram of the Gaia RC/RGB data (see Section 2) binned in $37.5$ pc cubes along slices in the $x-y$ plane at $z=-0.25$ (top), 00 (center), and $0.25$ (bottom) kpc. Center: MAF model of the observed number density $n_{\rm obs}$ of stars evaluated over the same slices in the $x-y$ plane at intervals of $12.5$ pc. Right: Dust-corrected model of the number density $n_{\rm corr}$ evaluated at the same points as $n_{\rm obs}$. In each row, the color scale is set to the $99^{\rm th}$ percentile value of the plotted $n_{\rm corr}$.
@@
This novel dust correction technique provides us with a much-improved PSD and solution to the CBE. In Eq. (1.3), the potential $\Phi$ enters at the same order as $f_{\rm corr}$, and so we expect an accurate $f_{\rm corr}$ to yield an accurate $\Phi$. However, downstream quantities such as the gravitational acceleration $\vec{a}$ and mass density $\rho$ enter at the same order as the first- and second-order gradients of $f_{\rm corr}$ respectively, and are therefore much more sensitive to mismodeling of $f_{\rm corr}$. Although analysis of $\vec{a}$ and $\rho$ are reserved for our companion paper Ref. [43], we now assess our dust-correction procedure in the most challenging region of space: in the disk towards the Galactic center.
@@
To examine the quality of our corrected phase space density in this challenging region of space, we show $n_{\rm corr}$ alongside $n_{\rm obs}$ in Figure 8 along tomographic slices in the $x-y$ plane in the disk. At $z=0$, exactly in the midplane, the expected structure of the disk emerges in $n_{\rm corr}$ from the heavily obscured $n_{\rm obs}$. Note that this is a purely data-driven dynamical inference of the structure of the Milky Way’s disk, as the $\epsilon$ network was not given any inductive bias regarding the structure of the Galactic disk. However, we can see in Figure 8 likely-spurious residual structures, both at small scales and large. Examples of the small-scale mismodeling include the “fringing” seen in $n_{\rm corr}$ at $z=0$ and $x\sim 5$ kpc, and an unexpected decrease in $n_{\rm corr}$ very close to the edge of the observation volume towards the Galactic center. At larger scales, there is a notable asymmetry in the overall number density of stars between $y>0$ and $y<0$ at small $x$. The fringing in particular is likely to affect the quality of the downstream inference for $\vec{a}$ and $\rho$, discussed in [43].
@@
Figure 9: $n_{\rm corr}$ (blue) and $n_{\rm obs}$ (gold) with their $1\sigma$ uncertainties evaluated along scan lines in Galactic cylindrical radius $R$ originating from $(x,y)=(0,0)$ in the directions $\phi=(10,0,-10)^{\circ}$ (left, center, and right), each at $z=(-0.25,0,0.25)$ kpc (top, center, bottom).
We note that the two-dimensional plots in Figure 8 reflect the mean output of the observed and corrected MAFs. This obscures the uncertainties inherent in our results. In Figure 9, we plot the $n_{\rm obs}$ and $n_{\rm corr}$ as a function of cylindrical radius $R$ along rays emanating from $(x,y)=(0,0)$ (at $\phi=-10^{\circ}$, $0^{\circ}$, $+10^{\circ}$, and $z=-0.25$, 00, and $+0.25$ kpc) including the full error budget as described in Appendix C. As can be seen, the errors in $n_{\rm corr}$ increase towards the Galactic center and within the disk, indicating the lack of confidence the $\epsilon$ network has in a particular solution as the data becomes sparser and regression of $\epsilon$ more difficult. This is the expected and desired behavior: similar uncertainties will propagate into results for $\Phi$ and downstream observables, avoiding overconfident predictions unsupported by the data. However, as an important exception to this, we do note that in some regions (particularly at $z=-0.25$ kpc) the decrease in $n_{\rm corr}$ towards the Galactic center remains statistically significant even as the errors increase. This behavior is unexpected, and may reflect modeling issues at the boundary of our data volume that have not been fully accounted for in the error budget.
@@ 5 Discussion
We have developed a novel dust correction technique for astrometric datasets of equilibrated or approximately-equilibrated stellar tracer populations. By simultaneously training neural networks that model the dust efficiency and Galactic potential using the CBE as a loss, we can estimate the underlying true PSD of stars in regions of the Galactic disk that are almost completely obscured by dust. In this paper, we described the algorithm and the training, along with investigating the resulting dust efficiency map. In our companion work [43], we show results for the $\Phi$ network, extracting a three-dimensional map of the gravitational potential, accelerations, and mass density across our $4$ kpc analysis volume.
Overall, the dust efficiency field $\epsilon(\vec{x})$ presented in this work is in excellent agreement with a recent high-resolution 3D dust map (Ref. [42]). However, some discrepancies and challenges remain, especially in the disk towards the Galactic center where there are many regions of heavy dust extinction with extremely sparse data. The $\epsilon$ network observed an increasing density along some of the few clear lines of sight towards the Galactic center, but it did not infer that the density should increase along other nearby lines of sight. This led to a decrease in $n_{\rm corr}$ below Galactic radii of $R<5$ kpc, as well as a non-axisymmetric structure in $n_{\rm corr}$ within $65<R<$ 6 kpc. However, we note that most of these unexpected features are not significant when uncertainties are considered.
@@
These challenges could also potentially be mitigated with future improvements to our technique. These include employing more advanced density estimation methods, such as diffusion models [48] or flow matching [49, 50], or learning $f$ using Fourier features [51] to directly address the small-scale spectral bias [52]. Another potential approach would involve modeling $f_{\rm corr}$ with a MAF and training $f_{\rm corr}$, $\epsilon$, and $\Phi$ in a single step by minimizing the CBE and the NLL simultaneously. Since $f_{\rm corr}$ is inherently smoother than $f_{\rm obs}$, this method would allow the MAF to focus less on modeling high-frequency dust-related features. While promising, these alternatives would require new hyperparameters and training regimens, which we leave for future work.
[…]

### v2

@@ 3.3 Regularization
This regularization term penalizes solutions where $\epsilon\neq 1$ over large portions of the sky and diverges as $\epsilon$ approaches 0, when $f_{\rm corr}(\vec{x})$ is no longer defined. This term introduces an additional hyperparameter $\lambda_{\epsilon}$. A hyperparameter scan over $\lambda_{\epsilon}$ reveals that small values of $\lambda_{\epsilon}$ fail to set a scale for $\epsilon$, leading to potentially inconsistent scales between different trainings of $\epsilon$ which will later be averaged together. Very large values of $\lambda_{\epsilon}$ force $\epsilon=1$ across the entire sky, a solution that fails to correct any of the effects of dust. Through numerical experiments, we found that the range $\lambda_{\epsilon}\in[10^{-2},1]$ yields stable results, and we ultimately settled on $\lambda_{\epsilon}=10^{-1}$ as the canonical value for this analysis.
@@ 4 Results
In this section, we present the results of the trained MAF model of $f_{\rm obs}$ as well as the $\epsilon$ network. In a companion paper Ref. [dustpaperII], we analyze the associated dust-corrected $\Phi$ model, where we examine the resulting acceleration and mass density field of the local Milky Way.
@@ 4.1 Dust Efficiency
    $\epsilon_{\rm L22}$ is computed from a Gaia dust map in the monochromatic band at $5500$ Å, whereas the RC/RGB sample of stars with 6D kinematic information were observed in the narrow $\sim 8600$ Å band of the Gaia RVS. Depending on their chemical composition, dust clouds will have slightly different morphology when observed in different bands.
@@
Figure 3: Left: CBE-derived $\epsilon$ projected in Galactic coordinates at distances $d=1$, 22, and $3$ kpc. At larger distances, familiar features in local dust clouds become more resolved in regions where $\epsilon<1$. However, not every feature seen here should be attributed to dust extinction. For example, the dark patch that first appears at $(\ell,b,d)\approx(5^{\circ},-4^{\circ},2\>\textrm{kpc})$ is an incomplete patch in the Gaia dataset due to crowding and saturation effects. Right: Same as left, using $\epsilon_{\rm L22}$ derived from the L22 dust map [Gaia2mass2022]. At 33 kpc, $\epsilon_{\rm L22}$ approaches 11 and is no longer reliable, as it cannot restore stars dimmed beyond the limits of the Gaia spectrometer.
@@
To understand the structure of $\epsilon$ both deep within the Milky Way’s disk as well as on the disk’s outskirts, we examine $\epsilon$ within tomographic slices through space at fixed $z$ in Figure 4. In the outer disk at $|z|=0.25$ kpc, minor dust lanes visible in $\epsilon_{\rm L22}$ are recovered in $\epsilon$. At larger radii, beyond the reach of our $\rm L22$} estimate of $\epsilon$, significant dust lanes near the Galactic center begin to appear. Within the center of the disk, shown in the central panel of Figure 4, we observe a complex structure of nearby dust clouds, which all mostly begin to appear within 0.520.5-2 kpc of the Earth. Interestingly, while most rays show monotonically decreasing $\epsilon$ with distance, select regions show $\epsilon$ slightly increasing with distance.
Figure 4: Left: Dust efficiency function $\epsilon$ estimated from the CBE along slices in the xyx-y plane at z=0.25z=-0.25 (top), $0.0$ (center), and $0.25$ kpc (bottom). The center row is a slice through the midplane of the Milky Way’s disk, where the most efficiency loss is expected due to the dense network of dust clouds. Right: $\epsilon_{\textrm{L22}}$ estimated from the extinctions provided by the L22 dust map, limited to within $d=2.5$ kpc (black circle). Beyond $d=2.5$ kpc, $\epsilon_{\rm L22}$ cannot restore missing stars and returns to $1$.
Figure 5: Comparison of $\epsilon_{\rm L22}$ derived from Ref. [Gaia2mass2022] (blue) and the $\epsilon$ estimated in this work (gold) along lines of sight centered on known significant dust clouds shown in Figure 1. The center line represents an average over a cone of opening angle $\theta=10$ arcmin. The error band for the $\epsilon$ model (blue) is dominated by measurement and statistical uncertainties. Likewise, the error band of the $\epsilon_{\rm L22}$ model (gold) includes the L22 dust map errors as well as an estimate of measurement and statistical errors. Both error bands include the subdominant variation across each circular section of the cone. The light gray band in some panels indicates lines of sight where $\epsilon_{\rm L22}$ degrades closer than $d=2.5$ kpc due to heavy dust extinction.
In Figure 5, we next examine $\epsilon$ evaluated within cones of radius $10$ arcmin that pass through the sixteen dusty regions of the Solar neighborhood highlighted in Figure 1. The uncertainties in $\epsilon$ shown in Figure 5 are derived by repeatedly retraining the MAFs and solving the CBE using multiple measurement error-smeared or resampled copies of the Gaia data, and by computing the variation across hundreds of retrainings with different network initializations. A full description of the uncertainty quantification procedure is given in Appendix C. The error band for $\epsilon_{\rm L22}$ is similarly derived, using the provided uncertainties in the L22 dust map [Gaia2mass2022] as well as error-smeared and resampled copies of the Gaia data. Additionally, for both $\epsilon$ and $\epsilon_{\rm L22}$, we take the average and variance over a short angular patch or “cone” in order to mitigate differences in resolution between the models. Variance across the cone is found to be subdominant to the primary measurement and statistical uncertainties.
@@
Figure 6: Pointwise “voxel” comparison between $\epsilon_{\rm L22}$ derived from Ref. [Gaia2mass2022] and the $\epsilon$ model that solves Eq. (1.3). Each efficiency value is evaluated over a grid of 3D voxels of volume $(25\>{\rm pc})^{3}$ that tile the space within $|z|<0.5$ kpc of and $d<2.5$ kpc. Top: Frequency of efficiency values in each voxel. Color scale indicates the density of voxels in each bin in a logarithmic scale. Bottom: Frequency of $\tau_{\epsilon}$ values between $\epsilon$ and $\epsilon_{\rm L22}$ accounting for errors. Dark blue indicates the $1-99\%$ percentile range of pull values. The solid black line indicates the average pull value.
As a final test, we compare $\epsilon$ and $\epsilon_{\rm L22}$ pointwise across a uniform, three-dimensional grid of nearby voxels in the Milky Way’s disk ($d<2.5$ kpc and $|z|<0.5$ kpc). In the left panel of Figure 6, we find that the models agree when there is no predicted extinction ($\epsilon\approx 1$). Additionally, we observe that $\epsilon$ lags behind as $\epsilon_{\rm L22}$ decreases from 1. This lag supports our previous observations that $\epsilon$ decreases from $1$ slightly later than $\epsilon_{\rm L22}$ along a given line of sight. This disagreement is slight, on the order of $\epsilon-\epsilon_{L22}||\approx 0.1$ when both $\epsilon$ and $\epsilon_{\rm L22}>0.4$. In practice, at a given $\vec{x}$ within a region of moderate dust extinction, this means that these models on average disagree at the 10% level on the fraction of missing stars. This is consistent with the level of disagreement seen in Figure 5. In the right panel of Figure 6, we present the pull $\tau_{\epsilon}=(\epsilon-\epsilon_{\rm L22})/\sqrt{\sigma_{\epsilon}^{2}+\sigma_{\epsilon_{\rm L22}}^{2}}$ between $\epsilon$ and $\epsilon_{\rm L22}$ evaluated at each voxel when accounting for uncertainties. As can be seen from the pull distribution, the bulk of $\tau_{\epsilon}$ values skew slightly positive, indicating that the amount of dust extinction tends to be slightly underestimated relative to the L22 dust map. However, the absolute value of this pull rarely exceeds $1\sigma$, indicating excellent agreement between these two models.
@@ 4.2 Dust-Corrected PSD
Figure 7: Left: MAF model of the observed number density $n_{\textrm{obs}}$ of Gaia RC/RGB stars projected in Galactic coordinates at distances $d=1$ (top), 22 (center), and 33 kpc (bottom). Right: Extinction-corrected number density $n_{\rm corr}\equiv n_{\textrm{obs}}/\epsilon$.
Figure 8: Left: Histogram of the Gaia RC/RGB data (see Section 2) binned in $37.5$ pc cubes along slices in the xyx-y plane at z=0.25z=-0.25 (top), 0 (center), and $0.25$ (bottom) kpc. Center: MAF model of the observed number density $n_{\rm obs}$ of stars evaluated over the same slices in the xyx-y plane at intervals of $12.5$ pc. Right: Dust-corrected model of the number density $n_{\rm corr}$ evaluated at the same points as $n_{\rm obs}$. In each row, the color scale is set to the $99^{\rm th}$ percentile value of the plotted $n_{\rm corr}$.
@@
This novel dust correction technique provides us with a much-improved PSD and solution to the CBE. In Eq. (1.3), the potential $\Phi$ enters at the same order as $f_{\rm corr}$, and so we expect an accurate $f_{\rm corr}$ to yield an accurate $\Phi$. However, downstream quantities such as the gravitational acceleration $\vec{a}$ and mass density $\rho$ enter at the same order as the first- and second-order gradients of $f_{\rm corr}$ respectively, and are therefore much more sensitive to mismodeling of $f_{\rm corr}$. Although analysis of $\vec{a}$ and $\rho$ are reserved for our companion paper Ref. [dustpaperII], we now assess our dust-correction procedure in the most challenging region of space: in the disk towards the Galactic center.
@@
To examine the quality of our corrected phase space density in this challenging region of space, we show $n_{\rm corr}$ alongside $n_{\rm obs}$ in Figure 8 along tomographic slices in the xyx-y plane in the disk. At $z=0$, exactly in the midplane, the expected structure of the disk emerges in $n_{\rm corr}$ from the heavily obscured $n_{\rm obs}$. Note that this is a purely data-driven dynamical inference of the structure of the Milky Way’s disk, as the $\epsilon$ network was not given any inductive bias regarding the structure of the Galactic disk. However, we can see in Figure 8 likely-spurious residual structures, both at small scales and large. Examples of the small-scale mismodeling include the “fringing” seen in $n_{\rm corr}$ at $z=0$ and $x\sim 5$ kpc, and an unexpected decrease in $n_{\rm corr}$ very close to the edge of the observation volume towards the Galactic center. At larger scales, there is a notable asymmetry in the overall number density of stars between $y>0$ and $y<0$ at small $x$. The fringing in particular is likely to affect the quality of the downstream inference for $\vec{a}$ and $\rho$, discussed in [dustpaperII].
@@
Figure 9: $n_{\rm corr}$ (blue) and $n_{\rm obs}$ (gold) with their $1\sigma$ uncertainties evaluated along scan lines in Galactic cylindrical radius $R$ originating from $(x,y)=(0,0)$ in the directions $\phi=(10,0,-10)^{\circ}$ (left, center, and right), each at z=(0.25,0,0.25)z=(-0.25,0,0.25) kpc (top, center, bottom).
We note that the two-dimensional plots in Figure 8 reflect the mean output of the observed and corrected MAFs. This obscures the uncertainties inherent in our results. In Figure 9, we plot the $n_{\rm obs}$ and $n_{\rm corr}$ as a function of cylindrical radius $R$ along rays emanating from $(x,y)=(0,0)$ (at $\phi=-10^{\circ}$, $0^{\circ}$, $10^{\circ}$, and z=0.25z=-0.25, 0, and +0.25+0.25 kpc) including the full error budget as described in Appendix C. As can be seen, the errors in $n_{\rm corr}$ increase towards the Galactic center and within the disk, indicating the lack of confidence the $\epsilon$ network has in a particular solution as the data becomes sparser and regression of $\epsilon$ more difficult. This is the expected and desired behavior: similar uncertainties will propagate into results for $\Phi$ and downstream observables, avoiding overconfident predictions unsupported by the data. However, as an important exception to this, we do note that in some regions (particularly at z=0.25z=-0.25 kpc) the decrease in $n_{\rm corr}$ towards the Galactic center remains statistically significant even as the errors increase. This behavior is unexpected, and may reflect modeling issues at the boundary of our data volume that have not been fully accounted for in the error budget.
@@ 5 Discussion
We have developed a novel dust correction technique for astrometric datasets of equilibrated or approximately-equilibrated stellar tracer populations. By simultaneously training neural networks that model the dust efficiency and Galactic potential using the CBE as a loss, we can estimate the underlying true PSD of stars in regions of the Galactic disk that are almost completely obscured by dust. In this paper, we described the algorithm and the training, along with investigating the resulting dust efficiency map. In our companion work [dustpaperII], we show results for the $\Phi$ network, extracting a three-dimensional map of the gravitational potential, accelerations, and mass density across our $4$ kpc analysis volume.
Overall, the dust efficiency field $\epsilon(\vec{x})$ presented in this work is in excellent agreement with a recent high-resolution 3D dust map (Ref. [Gaia2mass2022]). However, some discrepancies and challenges remain, especially in the disk towards the Galactic center where there are many regions of heavy dust extinction with extremely sparse data. The $\epsilon$ network observed an increasing density along some of the few clear lines of sight towards the Galactic center, but it did not infer that the density should increase along other nearby lines of sight. This led to a decrease in $n_{\rm corr}$ below Galactic radii of $R<5$ kpc, as well as a non-axisymmetric structure in $n_{\rm corr}$ within $5<R<6$ kpc. However, we note that most of these unexpected features are not significant when uncertainties are considered.
@@
These challenges could also potentially be mitigated with future improvements to our technique. These include employing more advanced density estimation methods, such as diffusion models [song2021scorebased] or flow matching [lipman2023flow, tong2024improving], or learning $f$ using Fourier features [10.5555/3495724.3496356] to directly address the small-scale spectral bias [pmlr-v97-rahaman19a]. Another potential approach would involve modeling $f_{\rm corr}$ with a MAF and training $f_{\rm corr}$, $\epsilon$, and $\Phi$ in a single step by minimizing the CBE and the NLL simultaneously. Since $f_{\rm corr}$ is inherently smoother than $f_{\rm obs}$, this method would allow the MAF to focus less on modeling high-frequency dust-related features. While promising, these alternatives would require new hyperparameters and training regimens, which we leave for future work.
Finally, we emphasize that the $\epsilon$ presented in this work serves a unique role compared to traditional dust maps. Instead of providing the magnitude of dust extinction, $\epsilon$ provides a direct estimate of the fraction of stars that dust extinction (or other effects, such as crowding) has removed from the sample. This serves as a direct estimate of the Gaia selection function in the RVS band. However, we caution that it is not possible to convert from $\epsilon$ to a traditional dust extinction map. Although we were able to convert the $\rm L22$} dust map into an effective $\epsilon_{\rm L22}$, this was only possible in limited regions of space where the Gaia survey was expected to be complete. Instead, we present $\epsilon$ as a parallel technique for mapping the effects of local dust clouds, not as a replacement for the available highly detailed 3D dust maps.